# MuLan: Adapting Multilingual Diffusion Models for Hundreds of Languages with Negligible Cost

### v1

@@ 4.1 Experimental setup
Datasets. We primarily use a subset of LAION-EN [36] with all samples that have aesthetic scores larger than 5.8 for training base models and PixArt [4] dataset for aesthetic models.
Implementation details. We trained MuLan adapters for a variety of pre-trained text encoders. By default, we use a transformer-based adapter with one encoder layer and one decoder layer. The number of transformer queries is set to 77 to match the input of Stable Diffusion [33].
All MuLan adapters are trained with AdamW [21] optimizer and 128 batch size. We use constant learning rate 1e-5 for Stable Diffusion 1.5/2.1 [33], 1e-6 for SDXL [28], and 2e-5 for Pixart-$\alpha$ [4]. For Stable Diffusion 1.5 [33], we train the adapter for 50k steps at the resolution of 512$\times$512, and for Stable Diffusion 2.1 [33] we adjust the resolution to 768$\times$768. For SDXL [28], we first train the adapter for 100k steps at the resolution of 512$\times$512 and finetune it for another 1k steps at the resolution of 1024$\times$1024. For Pixart-$\alpha$ [4], we train the adapter for 118k steps at the resolution of 512$\times$512. We randomly drop text conditions at the rate of 10% and use min-SNR [12] to accelerate training.
The training process was conducted on 8 NVIDIA A100-80G GPUs. For SD 1.5 and SD 2.1, training was completed within two days. For SDXL and Pixart-$\alpha$, training was completed within four days.
Evaluation Metrics. We use Crossmodal-3600 [41] to assess the model’s capabilities across 18 mainstream languages (denoted as XM18 hereafter). To evaluate the model’s generalization to additional languages, we tested multilingual versions of the COCO2014 [20] validation set and DPG-Bench [14]. We translated the prompts into 85 languages using Google Translate for these datasets. The model’s performance was compared with an ad-hoc translation-based SD1.5 model and other multilingual T2I models. Regarding evaluation metrics, we employed the Aesthetic Score to assess the quality of the generated images. Additionally, we calculated the Cosine Similarity (CLIP Sim) score with InternVL-LLaMA [6]. These metrics provide complementary insights into the visual quality and semantic accuracy of the model’s outputs.
@@ 4.2 Quantitive Results
Language InternVL-MuLan-SD15 Stable Diffusion (SD1.5)
AS CLIP Sim AS CLIP Sim
Arabic 6.33 38.22 6.11 34.43
English 6.35 38.61 6.06 39.57
French 6.29 37.96 6.05 38.16
German 6.24 37.95 6.01 38.37
Hindi 6.32 35.65 5.99 33.11
Italian 6.23 37.83 6.09 37.79
Japanese 6.27 37.96 6.02 36.67
Korean 6.33 37.57 6.06 36.36
Polish 6.32 37.07 6.09 36.69
Portuguese 6.32 37.04 6.08 37.19
Russian 6.33 37.89 6.02 37.38
Spanish 6.32 37.69 6.12 38.02
Thai 6.32 36.74 6.04 35.53
Turkish 6.35 37.97 6.09 37.68
Ukrainian 6.32 37.77 6.06 37.27
Vietnamese 6.25 37.94 6.04 37.20
Chinese 6.33 37.71 6.04 36.60
Dutch 6.36 37.41 6.12 37.59
Avg 6.31 37.61 6.06 36.97
Table 3: Comparison of zero-shot evaluation results between InternVL-Mulan-SD15 translation-based SD on XM18 [41]. InternVL-MuLan-SD15 consistently outperforms translation-based Stable Diffusion 1.5 in both aesthetic and CLIP similarity scores across most languages.
Language Model CLIP Sim
Chinese Taiyi-SDXL-3.5B [44] 35.44
Taiyi-SD-1B-Chinese [48] 36.84
@@
InternVL-Mulan-SDXL 37.21
Japanese JapaneseSDXL [39] 38
@@
Table 4: Comparison of CLIP similarity scores on Chinese and Japanese with specialized models on XM18. InternVL-MuLan-SD15 surpasses dedicated Chinese and Japanese models despite not using Chinese or Japanese data during training.
@@
Figure 3: Comparison of CLIP similarity scores of InternVL-MuLan-SD1.5 and AltDiffusion-m18 [46] across hundreds of languages on COCO2014 [20] val. Our model achieved performance comparable to AltDiffusion in mainstream languages while substantially surpassing AltDiffusion in less common languages.
Language Model Name Metrics
Overall Attribute Entity Global Other Relation
en SD1.5 62.66 74.02 72.95 72.11 74.95 75.28
AltDiffusion-m18 64.24 75.06 74.67 69.77 73.44 74.11
InternVL-MuLan-SD15 60.22 70.85 70.39 77.82 69.10 77.38
zh AltDiffusion-m18 62.18 72.33 73.58 71.22 72.55 74.30
InternVL-MuLan-SD15 58.75 71.04 71.33 66.44 71.74 70.23
ja AltDiffusion-m18 60.59 71.13 71.52 71.68 71.23 69.75
InternVL-MuLan-SD15 59.29 69.15 73.82 73.32 71.01 77.01
es AltDiffusion-m18 59.28 70.02 70.64 74.35 68.41 71.50
InternVL-MuLan-SD15 60.03 70.69 73.45 68.98 72.66 72.40
fr AltDiffusion-m18 59.41 69.81 69.51 72.59 69.27 70.38
InternVL-MuLan-SD15 60.83 72.74 71.99 73.11 71.06 75.34
Table 5: Performance metrics for different models across languages on DPG-Bench. InternVL-MuLan-SD15 achieves results close to those of AltDiffusion-m18 across various languages.
@@
Multilingual T2I Comparison. We integrated InternVL-LLaMA [6] into the adapter’s model, which we call InternVL-MuLan. Specifically, for each version of Stable Diffusion, we train a separate adapter model, which we refer to as InternVL-MuLan-SD15, InternVL-MuLan-SD21, and InternVL-MuLan-SDXL.
We first compared the performance of InternVL-MuLan-SD15 with translation-based Stable Diffusion 1.5 [33]. The evaluation results on XM18 [41] are shown in Table 3. Our model significantly outperforms SD1.5 in terms of image aesthetic quality. For CLIP similarity scores, our model surpasses Stable Diffusion in most languages.
@@
Next, we compared the performance of our model with AltDiffusion [46] in both mainstream and minority languages. We used DPG-Bench [14] to assess instruction-following ability in 5 mainstream languages. As shown in Table 5, our model’s instruction-following ability naturally extends to multiple languages, achieving results comparable to AltDiffusion [46]. To further evaluate our model’s capability in the broader range of minority languages, we evaluated our model on the COCO2014 [20] validation set (85 languages) and computed CLIP similarity scores, as shown in Figure 3(the complete list of scores will be provided in the appendix). Our model achieved performance comparable to AltDiffusion [46] in mainstream languages while substantially surpassing AltDiffusion [46] in less common languages.
Our model achieves state-of-the-art performance across multiple languages, reaching SOTA levels in mainstream languages and excelling in minority languages. Even though it was trained using only English text-image pairs, it outperforms certain specialized models in specific languages. While its performance in mainstream languages is slightly below that of AltDiffusion-m18 [46], the data and training costs we incurred are significantly lower than those of AltDiffusion. Specifically, our training on InternVL-MuLan-SD15 required only $0.5\times 8$ GPU-days, compared to the $19\times 64$ GPU-days needed by AltDiffusion-m18, highlighting a substantial difference in resource consumption. This makes our method both highly efficient and effective.
@@
Figure 4: Examples of our model integrates with community tools. MuLan can be equipped with MVDream [38] to generate 3D models, and it is also directly compatible with existing community tools such as LoRA [13], ControlNet [49], and LCM [23].
Adapter Type #params CLIP
MLP 30M
1-layer Transformer 18M 38.6
4-layer Transformer 67M 38.3
Table 6: Comparison of Different Adapter Types for InternVL-MuLan-SD15. The 4-layer transformer refers to a Transformer adapter with 4 encoder layers and 4 decoder layers. CLIP similarity scores are reported on XM18 [41]. ✗means the model generates images of extremely low quality.
Ablation Study on Adapter Architecture. We studied the impact of adapter types and the number of transformer layers on InternVL-LLaMA [6]. The results are shown in Table 6. The results show that the Transformer adapter outperforms the MLP adapter under different text encoder choices. The possible reason is that synonymous prompts have different word orders in different languages, and the Transformer adapter can possess the representational capabilities of a language model.
Comparison of Different Text Encoders. We trained Transformer adapters with frozen SD 1.5 on BERT [8], T5 [30], LLaMA [42, 9], and their image-centered multilingual alignment versions (if available).
Model Align Method en es fr zh
LLaMA2-7B [42]
mT5-xl [45]
MultiLang-CLIP [3] Language-centered 35.8 33.6 34.2 31.8
AltClip-m18 [5] Language-centered 35.2 33.9 34.2 31.9
ViT-XLM-R-L* [7] Image-centered 36.5 36.7 37.2 34.5
InternVL-LLaMA [6] Image-centered 38.6 37.7 38.0 37.7
Table 7: Text Encoder Clip Similarity Scores Across Different Languages. ViT-XLM-R-L* stands for CLIP-ViT-H-14-frozen-xlm-roberta-large-laion5B-s13B-b90k [16].
The results are shown in Table 7. We found that for multilingual text encoders trained purely on language, such as XLM-RoBERTa-large [7] and LLaMA [42, 9], the performance is poor when training an adapter with only English text-image pairs. They are unable to generate coherent images when we freeze these text encoders. Moreover, they cannot generalize to other languages through training on English text-image pairs. However, when we train these text encoders using the image-centered approach or align them to CLIP text encoder [29] using the language-centered method, MuLan can connect them to Stable Diffusion [33], enabling the T2I model to achieve multilingual image generation capability.

### v2

@@ 4.1 Experimental setup
Datasets. We primarily use a subset of LAION-EN (Schuhmann et al., 2022) with all samples that have aesthetic scores larger than 5.8 for training base models and PixArt (Chen et al., 2023a) dataset for aesthetic models.
Implementation details. We use InternVL-LLaMA (Chen et al., 2023b) as our text encoder, which has undergone Image-Centered Alignment on a multilingual image-text pair dataset. Then, we trained the MuLan Adapter for different image decoders using the architecture described in Section 3.
All MuLan adapters are trained with AdamW (Loshchilov & Hutter, 2017) optimizer and 128 batch size. We use constant learning rate 1e-5 for SD 1.5/2.1 (Rombach et al., 2022), 1e-6 for SDXL (Podell et al., 2023), and 2e-5 for Pixart-$\alpha$ (Chen et al., 2023a). For SD 1.5, we train the adapter for 50k steps at the resolution of 512$\times$512, and for SD 2.1 we adjust the resolution to 768$\times$768. For SDXL, we first train the adapter for 100k steps at the resolution of 512$\times$512 and finetune it for another 1k steps at the resolution of 1024$\times$1024. For Pixart-$\alpha$, we train the adapter for 118k steps at the resolution of 512$\times$512. We randomly drop text conditions at the rate of 10% and use min-SNR (Hang et al., 2023) to accelerate training.
The training process was conducted on 8 NVIDIA A100-80G GPUs. For SD 1.5 and SD 2.1, it took 12 hours, and for SDXL and Pixart-$\alpha$, it took two days.
Evaluation Metrics. We use XM3600 (Thapliyal et al., 2022) to assess the model’s capabilities across 12 mainstream languages (denoted as XM12 hereafter). To evaluate the model’s generalization to additional languages, we tested multilingual versions of the COCO2014 (Lin et al., 2015) validation set. We translated the prompts into 85 languages using Google Translate for these datasets. The model’s performance was compared with an ad-hoc translation-based SD 1.5 model and other multilingual T2I models (Yan et al., 2024). Regarding evaluation metrics, we employed FID and CLIP Score (SIM) calculated by InternVL-LLaMA (Chen et al., 2023b).
@@ 4.2 Quantitive Results
avg. en fr es it zh ja hi de ko ru ar pl
GlueGen* SIM 35.6 39.6 35.9 35.5 35.0 35.3 36.3
FID 17.4 21.4 17.1 18.5 17.6 17.6 16.1
AltDiffusion SIM 38.8 40.0 40.0 39.4 39.7 39.0 39.4 34.1 39.0 38.5 38.6 38.3 39.2
FID 9.1 8.1 8.2 9.4 10.4 9.4 9.4 10.2 8.2 8.9 9.7 8.9 9.1
SD15(NLLB) SIM 36.5 39.6 37.6 37.4 37.2 36.2 36.2 32.7 37.7 35.9 36.8 33.9 36.3
FID 25.0 21.4 24.9 25.0 25.0 25.5 24.9 28.2 24.5 25.4 24.6 26.6 24.4
SD15(Google) SIM 36.9 39.6 38.2 38.0 37.8 36.6 36.7 33.1 38.4 36.4 37.4 34.4 36.7
FID 21.6 21.4 21.6 23.9 22.4 19.1 20.5 19.6 22.1 21.7 21.6 22.9 21.8
PixArt(NLLB) SIM 38.3 39.0 39.7 39.4 39.2 38.1 38.6 34.4 40.3 37.7 39.3 35.6 38.7
FID 13.2 12.6 13.2 14.6 12.5 13.0 13.1 12.4 12.9 12.4 13.0 14.3 14.7
PixArt(Google) SIM 39.7 39.0 41.2 40.6 41.0 39.9 40.8 34.7 41.5 39.0 40.5 39.0 39.9
FID 13.7 12.6 13.7 15.5 12.9 13.8 13.9 12.3 12.9 13.2 12.9 15.1 15.4
SIM 37.7 38.6 38.0 37.7 37.8 37.7 38.0 35.6 38.0 37.6 37.9 38.2 37.1
Mulan-SD15 FID 14.4 13.9 14.7 16.5 13.5 17.0 13.7 13.0 13.8 12.8 13.8 14.0 16.5
SIM 39.5 39.6 40.2 39.6 40.0 39.3 40.5 37.2 40.5 39.3 39.6 39.1 39.1
Mulan-PixArt FID 11.8 12.8 12.2 13.6 11.7 10.6 10.7 13.4 12.4 8.4 10.5 11.9 13.2
Table 3: Comparison of Multilingual Generation Capabilities on XM12. GlueGen and AltDiffusion are other multilingual T2I methods, where the inputs for SD15 and PixArt-$\alpha$ are translated using Google Translate and NLLB. GlueGen is equivalent to the original SD15 when accepting English prompts. Its performance in English was not included in the average calculation.
lang Id Model SIM
zh Taiyi-SDXL-3.5B 35.44
Taiyi-SD-1B-Chinese 36.84
@@
ja JapaneseSDXL 38.00
@@
Table 4: Comparison of CLIP Score on Chinese and Japanese with specialized models on XM3600. InternVL-MuLan-SD15 surpasses specialized Chinese and Japanese models.
@@
Figure 3: Comparison of CLIP Score of InternVL-MuLan-SD15 and AltDiffusion across hundreds of languages on COCO2014 val. Our model achieved performance comparable to AltDiffusion in mainstream languages while substantially surpassing AltDiffusion in less common languages.
@@
Multilingual T2I Comparison. We integrated InternVL-LLaMA (Chen et al., 2023b) into the adapter’s model, which we call InternVL-MuLan. Specifically, for each image decoder, we train a separate adapter model, which we refer to as InternVL-MuLan-SD15, InternVL-MuLan-SD21, InternVL-MuLan-SDXL and InternVL-MuLan-PixArt.
As shown in Table 3, we compared our model with other multilingual image generation models. On the XM12 dataset, InternVL-MuLan-PixArt outperformed AltDiffusion (Ye et al., 2023a) and GlueGen (Qin et al., 2023) in CLIP score, highlighting its strong multilingual understanding. However, InternVL-MuLan-PixArt fell behind AltDiffusion in FID. To further evaluate image quality, we used Laion’s aesthetic predictor. InternVL-MuLan-SD15 and InternVL-MuLan-PixArt achieved average aesthetic scores of 6.31 and 6.67, respectively, surpassing AltDiffusion’s score of 6.05. This highlights the superior quality and visual appeal of images generated by our model.
@@
To further evaluate our model’s capability in the broader range of minority languages, we evaluated our model on the COCO2014 (Lin et al., 2015) validation set (85 languages) and computed CLIP Score, as shown in Figure 3. Our model achieved performance comparable to AltDiffusion in mainstream languages while substantially surpassing AltDiffusion in less common languages.
Our model achieves top performance across languages, reaching SOTA levels in mainstream languages and excelling in low-resource ones, all with minimal training costs. InternVL-MuLan-SD15 requires just 0.5$\times$8 GPU-days, and InternVL-MuLan-PixArt only 2$\times$8 GPU-days. In comparison, GlueGen costs 5 GPU-days yet performs worse than the low-cost InternVL-MuLan-SD15, while AltDiffusion demands at least 19$\times$64 GPU-days. This highlights the efficiency and cost-effectiveness of our approach.
Comparison of MuLan and Translate API We compared MuLan with the approach of using Google Translate and NLLB (NLLB Team et al., 2022) to convert non-English languages into English for input into image generation models (Sun et al., 2024). Non-English languages in XM12 were translated into English, as input of SD15 (Rombach et al., 2022) and PixArt-$\alpha$ (Chen et al., 2023a), with results compared to direct input into MuLan. As shown in Table 3, MuLan outperforms the approach that uses the open-source model NLLB as a translation tool and achieves performance comparable to using Google Translate. It demonstrates even better results in low-resource languages where translation tools often struggle to provide accurate translations.
Comparison of MuLan and “as-is” Baseline We evaluated Stable Diffusion v1.5 and PixArt-$\alpha$ by directly inputting prompts in 11 languages from the XM12 dataset, using InternVL-LLaMA to compute CLIP scores, shown in Table 5. These “as-is” results, where the prompts in different languages were fed into the T2I models without translation, were then compared with our MuLan-adapted models. We found that MuLan adapters consistently improved performance across all languages. Notably, even for languages close to English—such as French, Spanish, and German—our method still achieved clear gains over the as-is baseline, demonstrating its broad effectiveness.
Metric avg. en fr es it zh ja hi de ko ru ar pl
SD15(as-is) SIM 29.1 39.6 34.2 34.2 31.8 26.6 26.9 26.8 31.9 24.6 23.0 24.5 24.8
PixArt(as-is) SIM 29.8 39.0 38.2 36.9 36.0 24.2 22.1 24.0 36.8 22.4 28.9 22.2 27.0
Mulan-SD15 SIM 37.7 38.6 38.0 37.7 37.8 37.7 38.0 35.6 38.0 37.6 37.9 38.2 37.1
Mulan-PixArt SIM 39.5 39.6 40.2 39.6 40.0 39.3 40.5 37.2 40.5 39.3 39.6 39.1 39.1
GlueGen SIM(Laion) 21.1 22.3 21.3 20.6 19.6 21.5 21.0
AltDiffusion SIM(Laion) 23.1 24.2 23.6 23.2 23.1 24.3 23.0 21.1 24.3 22.5 22.3 22.7 23.2
SD15(Google) SIM(Laion) 22.3 22.3 23.8 23.4 22.8 22.6 21.3 20.8 24.5 21.7 22.6 19.5 21.9
PixArt(Google) SIM(Laion) 24.3 24.1 24.2 23.8 24.5 26.4 25.1 21.6 26.5 23.1 24.6 23.3 23.9
Mulan-SD15 SIM(Laion) 23.0 21.8 23.6 23.2 22.7 23.6 24.2 22.7 24.5 21.8 23.0 22.3 22.9
Mulan-PixArt SIM(Laion) 24.2 24.4 23.9 23.4 24.2 25.7 24.8 23.2 25.8 23.4 24.6 23.1 23.4
GlueGen VQAScore 0.53 0.81 0.51 0.52 0.45 0.41 0.50
AltDiffusion VQAScore 0.58 0.71 0.60 0.60 0.59 0.50 0.53 0.62 0.49 0.56 0.51 0.64 0.62
SD15(Google) VQAScore 0.61 0.81 0.66 0.71 0.61 0.57 0.50 0.67 0.50 0.70 0.68 0.41 0.52
PixArt(Google) VQAScore 0.75 0.85 0.78 0.80 0.82 0.71 0.71 0.82 0.71 0.76 0.76 0.59 0.69
Mulan-SD15 VQAScore 0.64 0.80 0.74 0.76 0.68 0.60 0.52 0.53 0.57 0.65 0.75 0.45 0.57
Mulan-PixArt VQAScore 0.74 0.88 0.81 0.83 0.81 0.71 0.71 0.61 0.73 0.74 0.79 0.59 0.72
Table 5: Multidimensional Comparison of Multilingual Image Generation on XM12. Part 1: Comparison between English-only T2I models using as-is prompts and MuLan; Part 2: Using the multilingual-capable Laion’s CLIP-ViT (LAION, 2023) for image-text similarity; Part 3: Evaluating fine-grained text-image alignment using VQAScore (Lin et al., 2024).
Alternative CLIP for Image-Text Similarity Image-text similarity is typically computed using the industry-standard CLIP-ViT models (Radford et al., 2021). However, these models are primarily trained on English data, making their similarity scores unreliable for non-English inputs. In this work, we employ InternVL-LLaMA, a multilingual model, to compute the image-text similarity scores. To further ensure the objectivity of the similarity evaluation, we additionally use the multilingual CLIP-ViT model released by Laion (LAION, 2023). The results shown in Table 5 demonstrate that our model remains robust, and the observed trends are consistent with those reported in Table 3.
Fine-grained Text-Image Alignment Evaluation While CLIP Score is useful for evaluating the presence of main subjects, it is limited in capturing object-level details and spatial relations—especially for compositional prompts in XM12. To address this, we adopt VQAScore (Lin et al., 2024) with GPT-4o as the evaluator, enabling reliable multilingual assessment. As shown in Table 5, our model achieves strong fine-grained alignment, outperforming GlueGen and AltDiffusion, and matching or surpassing translation-based baselines across most languages.
Comparison of MuLan and Specialized Models We compared the multilingual MuLan adapter with two language-specific models: Taiyi (Wu et al., 2024), trained on millions of Chinese image-text pairs, and the Japanese SDXL model (Shing & Akiba, ), fine-tuned on high-quality Japanese image-text pairs. Our model performs better on both languages, demonstrating its competitiveness with specialized models.
Comparison of Multilingual Semantic Alignment Methods In this study, we trained the MuLan Adapter on multilingual language models with and without Multilingual Semantic Alignment, using different alignment methods, and evaluated their performance on five non-English languages on the XM12 dataset. The results are shown in Table 6. We first selected the multilingual-supporting LLama (Touvron et al., 2023), mT5-xl (Xue et al., 2021) and XLM-R Large (Conneau, 2019) models as text encoders and trained MuLan without the Multilingual Semantic Alignment stage. Next, for Image-Centered Alignment, we choose InternVL-LLaMA (Chen et al., 2023b) and Mul-OpenCLIP (Conneau, 2019) as text encoders, both of which have be trained on multilingual image-text pairs such as Laion-5B (Schuhmann et al., 2022). For Language-Centered Alignment, we selected MultiLang-CLIP (Carlsson et al., 2022) and AltClip-m18 (Chen et al., 2022), models trained on parallel datasets like CCMatrix (Schwenk et al., 2020), We also trained an XLM-R Large on CCMatrix for comparison. Among them, Mul-OpenCLIP, MultiLang-CLIP, and AltClip-m18 all use the XLM-Roberta-Large architecture, but with different training methods or datasets. All image decoders used are based on SD 1.5 (Rombach et al., 2022).
Model Align Method avg SIM
LLaMA2-7B None
mT5-xl None
XLM-R Large None
MultiLang-CLIP LC 33.2
AltClip-m18 LC 33.3
XLM-R Large* LC 34.7
Mul-OpenCLIP IC 36.1
InternVL-LLaMA IC 37.8
Table 6: CLIP Score on five non-English languages of Different Text Encoders trained with MuLan Adapter. Mul-OpenCLIP stands for CLIP-ViT-H-14-frozen-xlm-roberta-large-laion5B-s13B-b90k (LAION, 2023). “LC” and “IC” represent Language-Centered Alignment and Image-Centered Alignment, respectively; XLM-R Large* is trained by the authors using CCMatrix.
The results are shown in Table 6. We found that multilingual text encoders trained solely on multilingual corpora, such as LLaMA, mT5-xl, and XLM-R Large, fail to understand non-English prompts entirely. This is because these models have not undergone multilingual alignment, so their multilingual embeddings are not aligned. After connecting these language models with the MuLan adapter and training with English text-image pairs, only the English embeddings are aligned with the visual feature space, preventing understanding of non-English prompts. However, with either of the two alignment methods, the multilingual embeddings are first mapped to a unified feature space, which the MuLan adapter then aligns with the image space using only English image-text pairs, enabling multilingual image generation.
@@
Figure 4: Examples of MuLan integrates with community tools. MuLan seamlessly integrates with MVDream (Shi et al., 2023) for 3D model generation and is fully compatible with community tools like LoRA (Hu et al., 2021), ControlNet (Zhang et al., 2023), and LCM (Luo et al., 2023).
The results also suggest that Image-Centered Alignment outperforms Language-Centered Alignment by a significant margin. We speculate that this is because, in Image-Centered Alignment, language alignment is achieved through the image, which acts as a bridge between the language embeddings and the visual space. As a result, the semantic space of different languages is already aligned with the image prior to training with MuLan, allowing MuLan to further refine this alignment. In contrast, Language-Centered Alignment relies entirely on MuLan to establish the connection between language embeddings and the image feature space, making the MuLan adapter training more challenging and resulting in lower performance.
In summary, MuLan innovatively uses an image-centered aligned language model for multilingual image generation. Previous methods like GlueGen (Qin et al., 2023) and AltDiffusion (Ye et al., 2023a), which rely on language-centered alignment, face inherent limitations due to their alignment method. GlueGen aligns non-English embeddings with the CLIP text encoder used in SD, with SD kept frozen, but as shown in Table 3, its performance on non-English languages is significantly worse than that on English. AltDiffusion, by contrast, retrains both text encoder and UNet on billions of multilingual pairs, incurring high training costs. In comparison, our approach achieves strong, balanced multilingual generation with minimal cost, requiring only a small adapter while keeping the SD frozen. In practice, the use of a unified, image-centered alignment pre-trained model allows easy adaptation to various SD models in the community with minimal overhead.